OpenStudy (anonymous):

Convert 149° in Fahrenheit to Celsius temperature using the formula . f=9/5c+32lol 75°C 70°C 65°C 60°C

OpenStudy (angela210793):

what formula? O.o

OpenStudy (anonymous):

i got b

OpenStudy (ghazi):

60

OpenStudy (ganpat):

f = (9/5)C + 32 149 = (9/5) C + 32 (9/5)C = 117 C = 65 degree c .. option c..